OpenStudy (anonymous):

If y= -252 when x = 84, what is y when x = 74?

OpenStudy (gorv):

\[\frac{ y }{ x }=\frac{ -252 }{ 84}\]

OpenStudy (gorv):

y/x =-3

OpenStudy (gorv):

multiply by x on both side y=-3x at x=74 y=-3*74=-224

OpenStudy (anonymous):

I got -222?